OpenStudy (anonymous):

The ratio of 8 less than a number to 28 more than that number is 2 to 5

OpenStudy (dan815):

(x-8):(x+28) = 2:5

OpenStudy (dan815):

like this so

OpenStudy (anonymous):

Let x represent the number,\[\begin{align}x-8 : 28 + x &= 2 : 5 \\ \frac{x - 8}{28+x} &= \frac{2}{5} \end{align}\] Cross multiply and solve for x.

OpenStudy (dan815):

(x-8)/((x-8+x+28) = 2/(5+2) (x-8)/(2x+20)=2/7 7(x-8)=2(2x+20) 7x-56=4x+40 3x=96 x=96/3 =32

OpenStudy (anonymous):

i forgot to add something to it what is the number is it 32,30,28,15 ??

OpenStudy (dan815):

I hope you get how to do this on your own now.

OpenStudy (anonymous):

damn haha xp thank you :)